OpenStudy (anonymous):

\[|x-9|=16~~~~~~~~~~~~solve~~~~~~~~~~\]

OpenStudy (anonymous):

x-9 = 16 and x-9 = -16 so, x= 25 and x = -7

OpenStudy (campbell_st):

there are 2 equations, the easy way is to allow the RHS to be positive and negative x - 9 = 16 and x - 9 = -16 solve the 2 equations for the solutions.

OpenStudy (anonymous):

Um, what of |x^2 - 9| =16?

OpenStudy (anonymous):

other way is to solve it by graph.